OpenStudy (anonymous):

find dy/dx by implicit differentiation. tan(x-y)=y/(7+x^(2))

zepdrix (zepdrix):

Oh boy this one is going to be a little messy. I dunno about you, but I really dislike using the quotient rule, so let's multiply that denominator to the other side.\[\large \left(7+x^2\right)\tan(x-y)=y\] From here, we'll take the derivative implicitly, applying the product rule on the left.

zepdrix (zepdrix):

\[\large \color{royalblue}{\left(7+x^2\right)'}\tan(x-y)+\left(7+x^2\right)\color{royalblue}{\tan(x-y)'}=\color{royalblue}{y'}\] So here is the setup for the product rule, we'll have to take the derivative of eveyrthing in blue. Understand how I set that up? :O

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I got 2x*tan(x-y)+(7+x^2)*sec^2(x-y)=y'

zepdrix (zepdrix):

Ok good, you just missed one simple step. You forget to apply the `Chain Rule` to the inside of the tangent.

OpenStudy (anonymous):

oh, so would it be 2x*tan(x-y)*y'+(7+x^2)*sec^2(x-y)=y'?

zepdrix (zepdrix):

Woops wrong tangent, the chain rule will be applied to the tangent that you took a derivative of, look at this to help maybe. \[\large (2x)\tan(x-y)+\left(7+x^2\right)\sec^2(x-y)\color{royalblue}{(x-y)'}=y'\]

OpenStudy (anonymous):

do you multiply by y' in the product rule when you take the derivative of something with a y term?

zepdrix (zepdrix):

Yes, kind of. The problem is, the inside of our tangent function wasn't simply `y`. You're probably used to dealing with problems like \(\large \tan y\) , yes in that case a y' would pop out when you take it's derivative. But this is due to the chain rule. You'll only have a y' when you take the derivative of y directly. I dunno, maybe that's a bad way to explain it. \[\large \frac{d}{dx}\tan y \qquad = \qquad \sec^2y(y)'\]The chain rule tells us to multiply by the derivative of the inside, so we'll do that,\[\large = \qquad (\sec^2y)y'\] ---------------------------- Our problem is a little trickier.\[\large \frac{d}{dx}\tan(x-y) \qquad =\qquad \sec^2(x-y)(x-y)'\]The derivative of the inside gives us,\[\large =\qquad \sec^2(x-y)(1-y')\]

OpenStudy (anonymous):

so then the product rule would be 2x*tan(x-y)+(7+x^2)*sec^2(x-y)*(1-y')=y'?

zepdrix (zepdrix):

After applying the product rule, we apply the chain rule to the second term, and yes that's what we get. Now comes a bit of tricky algebra. Moving stuff around, shuffle shuffle, try to solve get the y' alone on one side.

OpenStudy (anonymous):

that's where I get lost with implicit differentiation.

OpenStudy (anonymous):

I know you get the y' terms on one side together

zepdrix (zepdrix):

Do you understand how to get the y' on one side? Or need to see some steps? c:

OpenStudy (anonymous):

I need to see some steps.. not good with algebra..

zepdrix (zepdrix):

\[\large (2x)\tan(x-y)+\color{orangered}{\left(7+x^2\right)\sec^2(x-y)}(1-y')=y'\] Let's start by distributing the big... orange mess of stuff... into that brackets containing y' \[\large (2x)\tan(x-y)+\color{orangered}{\left(7+x^2\right)\sec^2(x-y)}-\color{orangered}{\left(7+x^2\right)\sec^2(x-y)}y'=y'\]

zepdrix (zepdrix):

We can then move the y' term over to the right side by adding it to both sides,\[\large (2x)\tan(x-y)+\left(7+x^2\right)\sec^2(x-y)=y'+\left(7+x^2\right)\sec^2(x-y)y'\]

zepdrix (zepdrix):

Then factor a y' ouch of each term.\[\large (2x)\tan(x-y)+\left(7+x^2\right)\sec^2(x-y)=y'\left[1+\left(7+x^2\right)\sec^2(x-y)\right]\]

zepdrix (zepdrix):

Then divide by the big mess...\[\large \frac{(2x)\tan(x-y)+\left(7+x^2\right)\sec^2(x-y)}{\left[1+\left(7+x^2\right)\sec^2(x-y)\right]}=y'\]

zepdrix (zepdrix):

We end up with something really awful, like that.... :\

OpenStudy (anonymous):

Thank you for showing what you did AND explaining the step because that's helpful for me. And thank you for taking the time to help! Also, the one thing I am unsure of is the distributing step.

zepdrix (zepdrix):

We took our giant ugly potato and distributed it (multiplied it by each term) inside the brackets. As illustrated here. What makes it confusing maybe is that it's a really big potato. So it's hard to understand that we're distributing that entire thing to each term. It almost looked like we made a copy of it, which maybe looked weird :)